# Simple Guidance Mechanisms for Discrete Diffusion Models

### v1

@@ 1 Introduction
However, applying controllable diffusion-based generation to tasks where the data is discrete (e.g., molecule design or text generation) presents challenges. First, standard diffusion models and their guidance mechanisms are not directly applicable, since they require taking gradients with respect to the data, and these are not defined in discrete settings. Second, popular discrete extensions of diffusion (Lou et al., 2023; Sahoo et al., 2024a) cannot perform multiple editing passes on generated tokens, hence are not ideal for controllable generation. Third, the performance of discrete diffusion models (measured by perplexity) lags behind autoregressive (AR) models, especially for classes of diffusion that are amenable to control, such as uniform noise (Austin et al., 2021; Lou et al., 2023).
Here, we propose discrete diffusion models and guidance mechanisms that are effective at controllable generation and that address the above challenges. First, we provide straightforward and easy-to-implement adaptations of classifier-based and classifier-free guidance for discrete diffusion models. Second, we revisit uniform noise diffusion language models (UDLM), which undo random token perturbations and are particularly amenable to guidance, since they can continuously edit discrete data (Austin et al., 2021). We address performance issues that plagued previous iterations of uniform noise discrete diffusion by deriving a continuous-time version of the evidence lower bound, which tightens the variational gap (Kingma et al., 2021; Sahoo et al., 2024a).
We demonstrate the effectiveness of guidance with discrete diffusion models on several domains: genomics, molecular generation, and discretized images. We find that discrete diffusion models are more controllable than autoregressive models when paired with classifier-free guidance. Moreover, our proposed classifier-based method improves upon previous guidance mechanisms for diffusion (Gruver et al., 2024) and AR models (Dathathri et al., 2019; Yang & Klein, 2021). Our language modeling experiments also reveal that contrary to a widely-held belief (Austin et al., 2021; Lou et al., 2023), uniform noise diffusion can attain state-of-the-art performance on small vocabulary datasets (e.g,. molecules, DNA) and that UDLM attains a new state-of-the-art in perplexity among uniform noise diffusion models.
@@ 2 Background
Notation  Let $\mathcal{V}$ be the space of all one-hot tokens over some vocabulary consisting of $N$ unique characters: $\mathcal{V}=\{{\mathbf{z}}\in\{0,1\}^{N}:\sum_{i}{\mathbf{z}}_{i}=1\}\subset\Delta^{N}$, where $\Delta^{N}$ represents the simplex over $N$ categories. Let ${\bm{1}}$ be a $N$-dimensional column vector of all ones, and denote the Hadamard product between two vectors as $\odot$.⊙ . We define ${\mathbf{z}}^{(1:L)}$ as a sequence of $L$ tokens, where ${\mathbf{z}}^{(\ell)}\in\mathcal{V},$ for all tokens $\ell\in 1,\ldots,L,$ and use $\mathcal{V}^{L}$ to denote the set of all such sequences. Finally, let $\mathrm{Cat}(\cdot;p)$ denote the categorical distribution with probability vector $p\in\Delta^{N}.$
@@ 2.1 Discrete Diffusion Models
Diffusion models are a class of generative models defined by a denoising network $p_{\theta}$ that is trained to remove noise from latent variables ${\mathbf{z}}_{t}$. These latents are generated by a fixed corruption process $q$ that, starting from clean data ${\mathbf{x}}_{0}$ drawn from the data distribution $q({\mathbf{x}}_{0})$, increasingly adds more noise to ${\mathbf{z}}_{t}$, as $t$ increases (Sohl-Dickstein et al., 2015; Song & Ermon, 2019; Ho et al., 2020).
In discrete denoising diffusion probabilistic models (D3PM; Austin et al. (2021)), the noising process is defined in terms of a transition matrix $Q_{t|s}$ whose $(i,j)^{\text{th}}$ entry is the probability of transitioning from the $i$-th state at time $s$ to the $j$-th state at time $t$. This induces a Markov corruption process where we have $q({\mathbf{z}}_{t}\mid{\mathbf{z}}_{s})=\mathrm{Cat}({\mathbf{z}}_{t};Q_{t|s}{\mathbf{z}}_{s})$.
@@
$q({\mathbf{z}}_{t}|{\mathbf{x}}_{0})=\mathrm{Cat}({\mathbf{z}}_{t};\alpha_{t}{\mathbf{x}}_{0}+(1-\alpha_{t})\bm{\pi}{}),$ (1)
@@
$q({\mathbf{z}}_{s}|{\mathbf{z}}_{t},{\mathbf{x}}_{0})=\mathrm{Cat}\left({\mathbf{z}}_{s};\frac{[\alpha_{t|s}{\mathbf{z}}_{t}+(1-\alpha_{t|s}){\bm{1}}\bm{\pi}^{\top}{\mathbf{z}}_{t}]\odot[\alpha_{s}{\mathbf{x}}_{0}+(1-\alpha_{s})\bm{\pi}]}{\alpha_{t}{\mathbf{z}}_{t}^{\top}{\mathbf{x}}_{0}+(1-\alpha_{t}){\mathbf{z}}_{t}^{\top}\bm{\pi}}\right).$ (2)
@@ 2.2 Diffusion Guidance
$\nabla_{{\mathbf{z}}_{s}}\log p^{\gamma}({\mathbf{z}}_{s}\mid y,{\mathbf{z}}_{t})$ $=\gamma\cdot[\nabla_{{\mathbf{z}}_{s}}\log p_{\theta}({\mathbf{z}}_{s}\mid y,{\mathbf{z}}_{t})-\nabla_{\mathbf{x}}\log p_{\theta}({\mathbf{z}}_{s}\mid{\mathbf{z}}_{t})]+\nabla_{{\mathbf{z}}_{s}}\log p_{\theta}({\mathbf{z}}_{s}\mid{\mathbf{z}}_{t})$
@@
where $p_{\theta}({\mathbf{z}}_{s}\mid y,{\mathbf{z}}_{t})$ and $p_{\theta}({\mathbf{z}}_{s}\mid{\mathbf{z}}_{t})$ represent conditionally and unconditionally trained diffusion models, respectively. In practice, this is often implemented by using a single model $p_{\theta}$ and randomly dropping out / masking the conditioner $y$ during training.
The problem with applying guidance to discrete diffusion is that guidance terms are not differentiable with respect to the the discrete representations ${\mathbf{z}}_{t}$.
@@ 3 Guidance Algorithms for Discrete Diffusion
We introduce two guidance algorithms for discrete diffusion that circumvent the issue of non-differentiability. As in Section 2.2, we formalize the guidance term as a probability $p(y|{\mathbf{z}})\in\Delta^{K}$, where $y\in\{1,\ldots,K\}$ is one of $K$ possible classes, and we begin with the distribution that controls the strength of the guidance term via the $\gamma$ temperature parameter:
$\log p^{\gamma}({\mathbf{z}}_{s}\mid y,{\mathbf{z}}_{t})=\gamma\log p(y\mid{\mathbf{z}}_{s},{\mathbf{z}}_{t})+\log p({\mathbf{z}}_{s}\mid{\mathbf{z}}_{t})+c,$ (5)
where $c$ denotes the logarithm of the normalization constant.
@@ 3.2 Classifier-based Guidance
Discrete Classifier-Based Guidance  We first introduce the following additional notation: given ${\mathbf{z}}^{1:L}$, let $\tilde{\mathcal{Z}}_{\ell}({\mathbf{z}}^{1:L})$ be the set of sequences $\tilde{{\mathbf{z}}}^{(1:L)}$ for which $\tilde{{\mathbf{z}}}^{(\ell^{\prime})}={\mathbf{z}}^{(\ell^{\prime})}$ for all $\ell^{\prime}\neq\ell,$ i.e., the set of sequences that are either the same as or only differ in position $\ell$ relative to ${\mathbf{z}}^{(1:L)}.$ We can sample from $p^{\gamma}({\mathbf{z}}_{s}^{(\ell)}\mid{\mathbf{z}}_{t}^{(1:L)},y)$ by training a classifier $p_{\phi}:\mathcal{V}^{L}\rightarrow\Delta^{K}$ on noised latents ${\mathbf{z}}_{t}^{(1:L)}$ for $t\in[0,1]$ and use this to model the first term on the right-hand side of (10) by only evaluating $p_{\phi}$ on sequences for which ${\mathbf{z}}_{s}^{(1:L)}$ and ${\mathbf{z}}_{t}^{(1:L)}$ differ by at most the token at position $\ell$:
@@
First Order Approximation  While tractable, this formulation suffers from the drawback that at each denoising step we must perform ${\mathcal{O}}(L\cdot N)$ forward passes through the classifier model, which can quickly become impractical for larger vocabularies. Similarly to Grathwohl et al. (2021), Vignac et al. (2022), we can treat the classifier $p_{\phi}$ as a continuous function of the one-hot inputs $\tilde{{\mathbf{z}}}^{(1:L)}\in\mathbb{R}^{L\times N}$ and use the first order Taylor approximation of $\log p_{\phi}$ to efficiently compute $p_{\phi}(y\mid\tilde{{\mathbf{z}}}^{(1:L)})$ with only a single forward and backward pass through the classifier model:
@@ 4 Uniform Diffusion Language Models
While masked diffusion models have demonstrated better language modeling compared to other discrete diffusion (Austin et al., 2021; Lou et al., 2023), we argue that they are less amenable to guidance, since once a token is unmasked at some time $t$ it remains so for all $s<t$. In contrast, for uniform noising, intermediate latents can be refined multiple times during the denoising process. We therefore revisit categorical uniform noise discrete diffusion, where $\bm{\pi}=\bm{u}:={\bm{1}}/N.$ Our aim is that by analyzing this class of diffusion models more carefully, we can reduce the gap to absorbing-state and yield performant models that are more easily steered by guidance tools we developed above.
@@ 4.1 Uniform Noise Diffusion
Discrete Time Likelihood Bound  We start with the variational lower bound that is defined by a general diffusion process. For discrete-time diffusion, i.e., some finite steps $T$, we define $t(i)=(i+1)/T$ and $s(i)=i/T$ for $i$ in $0,\ldots,T.$ The denoising network $p_{\theta}$ is trained to minimize a variational upper bound (NELBO), which is given by:
$\mathbb{E}_{q}\Bigg{[}\underbrace{-\log p_{\theta}({\mathbf{x}}_{0}|{\mathbf{z}}_{t(0)})}_{\begin{array}[]{c}{\mathcal{L}_{\text{recons}}}\end{array}}+\underbrace{\sum_{i=1}^{T}D_{\mathrm{KL}}[q({\mathbf{z}}_{s(i)}|{\mathbf{z}}_{t(i)},{\mathbf{x}}_{0})\|p_{\theta}({\mathbf{z}}_{s(i)}|{\mathbf{z}}_{t(i)})]}_{\begin{array}[]{c}{\mathcal{L}_{\text{diffusion}}}\end{array}}\Bigg{]}+\underbrace{D_{\mathrm{KL}}[q({\mathbf{z}}_{t(T)}|{\mathbf{x}}_{0})\|p_{\theta}({\mathbf{z}}_{t(T)})]}_{\begin{array}[]{c}{\mathcal{L}_{\text{prior}}}\end{array}},$ (13)
where $D_{\mathrm{KL}}$ represents the Kullback-Leibler divergence. When clear, we drop the explicit dependence of $t$ and $s$ on the discrete step $i$.
@@
$q({\mathbf{z}}_{s}\mid{\mathbf{z}}_{t},{\mathbf{x}}_{0})=\mathrm{Cat}\left({\mathbf{z}}_{s};\frac{N\alpha_{t}{\mathbf{z}}_{t}\odot{\mathbf{x}}_{0}+(\alpha_{t|s}-\alpha_{t}){\mathbf{z}}_{t}+(\alpha_{s}-\alpha_{t}){\mathbf{x}}_{0}+\frac{(\alpha_{s}-\alpha_{t})(1-\alpha_{s})}{N\alpha_{s}}{\bm{1}}}{N\alpha_{t}\langle{\mathbf{z}}_{t},{\mathbf{x}}_{0}\rangle+1-\alpha_{t}}\right)$ (14)
Denoising Process  The optimal form for the reverse diffusion process $p_{\theta}$ matches (14): in fact setting $p_{\theta}$ to (14) reduces the KL terms in (13) to zero. However, setting $p_{\theta}$ to exactly (14) is not possible because it cannot be a function ${\mathbf{x}}_{0}$ (which $p_{\theta}$ is generating). Therefore, we introduce a predictive model ${\mathbf{x}}_{\theta}({\mathbf{z}}_{t},t):\mathcal{V}\times[0,1]\rightarrow\Delta^{N}$ of the ‘clean’ data given a noisy latent ${\mathbf{z}}_{t}$ at time $t$. We use ${\mathbf{x}}_{\theta}$ to parameterize the denoising process as $p_{\theta}({\mathbf{z}}_{s}\mid{\mathbf{z}}_{t})=q({\mathbf{z}}_{s}\mid{\mathbf{z}}_{t},{\mathbf{x}}={\mathbf{x}}_{\theta})$, yielding:
@@
Note that this minimizes the ${\mathcal{L}_{\text{diffusion}}}$ term in (13) precisely when ${\mathbf{x}}_{\theta}={\mathbf{x}}_{0},$ as desired. To simplify notation, we omit below the explicit dependence of ${\mathbf{x}}_{\theta}$ on $t$.

### v3

@@ 1 Introduction
However, applying controllable diffusion-based generation to tasks where the data is discrete (e.g., molecule design or text generation) presents challenges. First, standard diffusion models and their guidance mechanisms are not directly applicable, since they require taking gradients with respect to the data, and these are not defined in discrete settings. Second, popular discrete extensions of diffusion (Sahoo et al., 2024a; Shi et al., 2024) cannot perform multiple editing passes on generated tokens, hence are not ideal for controllable generation. Third, the performance of discrete diffusion models (measured by perplexity) lags behind autoregressive (AR) models, especially for classes of diffusion that are amenable to control, such as uniform noise (Austin et al., 2021; Lou et al., 2023).
Here, we propose discrete diffusion models and guidance mechanisms that are effective at controllable generation and that address the above challenges. First, we provide straightforward and easy-to-implement adaptations of classifier-based and classifier-free guidance for discrete diffusion models. Second, we revisit uniform noise diffusion language models (UDLM), which undo random token perturbations and are particularly amenable to guidance, since they can repeatedly edit their samples (Austin et al., 2021) and thus correct errors. We address performance issues that plagued previous iterations of uniform noise discrete diffusion by deriving a continuous-time version of the evidence lower bound, which tightens the variational gap (Kingma et al., 2021; Sahoo et al., 2024a).
We demonstrate the effectiveness of guidance with discrete diffusion models on several domains: genomics, molecule strings, and discretized images. We find that discrete diffusion models are more controllable than AR when paired with classifier-free guidance. Moreover, our proposed classifier-based method improves upon previous guidance mechanisms for diffusion (Gruver et al., 2024) and AR (Dathathri et al., 2019; Yang & Klein, 2021). Our language modeling experiments also reveal that contrary to a widely-held belief (Austin et al., 2021; Lou et al., 2023), uniform noise diffusion can attain state-of-the-art performance on small vocabulary datasets (e.g,. molecules, DNA) and that UDLM attains a new state-of-the-art in perplexity among uniform noise diffusion models.
@@ 2 Background
Notation  Let $\mathcal{V}$ be the space of all one-hot tokens over some vocabulary consisting of $N$ unique characters: $\mathcal{V}=\{{\mathbf{z}}\in\{0,1\}^{N}:\sum_{i}{\mathbf{z}}_{i}=1\}\subset\Delta^{N}$, where $\Delta^{N}$ represents the simplex over $N$ categories. Let ${\bm{1}}$ be a $N$-dimensional column vector of all ones, and denote the Hadamard product between two vectors as $\odot$.⊙ . We define ${\mathbf{z}}^{(1:L)}$ as a sequence of $L$ tokens, where ${\mathbf{z}}^{(\ell)}\in\mathcal{V},$ for all tokens $\ell\in\{1,\ldots,L\},$ and use $\mathcal{V}^{L}$ to denote the set of all such sequences. Finally, let $\mathrm{Cat}(\cdot;p)$ denote the categorical distribution with probability vector $p\in\Delta^{N}.$
@@ 2.1 Discrete Diffusion Models
Diffusion models are a class of generative models defined by a denoising network $p_{\theta}$ that is trained to remove noise from latent variables ${\mathbf{z}}_{t}$. These latents are generated by a fixed corruption process $q$ that, starting from clean data ${\mathbf{x}}$ drawn from the data distribution $q({\mathbf{x}})$, increasingly adds more noise to ${\mathbf{z}}_{t}$, as $t$ increases (Sohl-Dickstein et al., 2015; Song & Ermon, 2019; Ho et al., 2020).
In discrete denoising diffusion probabilistic models (D3PM; Austin et al. (2021)), the noising process is defined in terms of a transition matrix $Q_{t|s}$ whose $(i,j)^{\text{th}}$ entry is the probability of transitioning to the $i$-th state at time $t$ given the $j$-th state at time $s$. This induces a Markov corruption process where we have $q({\mathbf{z}}_{t}\mid{\mathbf{z}}_{s})=\mathrm{Cat}({\mathbf{z}}_{t};Q_{t|s}{\mathbf{z}}_{s})$.
@@
$q({\mathbf{z}}_{t}|{\mathbf{x}})=\mathrm{Cat}({\mathbf{z}}_{t};\alpha_{t}{\mathbf{x}}+(1-\alpha_{t})\bm{\pi}{}),$ (1)
@@
$q({\mathbf{z}}_{s}|{\mathbf{z}}_{t},{\mathbf{x}})=\mathrm{Cat}\left({\mathbf{z}}_{s};\frac{[\alpha_{t|s}{\mathbf{z}}_{t}+(1-\alpha_{t|s}){\bm{1}}\bm{\pi}^{\top}{\mathbf{z}}_{t}]\odot[\alpha_{s}{\mathbf{x}}+(1-\alpha_{s})\bm{\pi}]}{\alpha_{t}{\mathbf{z}}_{t}^{\top}{\mathbf{x}}+(1-\alpha_{t}){\mathbf{z}}_{t}^{\top}\bm{\pi}}\right).$ (2)
@@ 2.2 Diffusion Guidance
$\nabla_{{\mathbf{z}}_{s}}\log p^{\gamma}({\mathbf{z}}_{s}\mid y,{\mathbf{z}}_{t})$ $=\gamma\cdot[\nabla_{{\mathbf{z}}_{s}}\log p_{\theta}({\mathbf{z}}_{s}\mid y,{\mathbf{z}}_{t})-\nabla_{{\mathbf{z}}_{s}}\log p_{\theta}({\mathbf{z}}_{s}\mid{\mathbf{z}}_{t})]+\nabla_{{\mathbf{z}}_{s}}\log p_{\theta}({\mathbf{z}}_{s}\mid{\mathbf{z}}_{t})$
@@
with $p_{\theta}({\mathbf{z}}_{s}\mid y,{\mathbf{z}}_{t})$ conditionally and $p_{\theta}({\mathbf{z}}_{s}\mid{\mathbf{z}}_{t})$ unconditionally trained diffusion models.
@@ 3 Guidance Algorithms for Discrete Diffusion
Applying guidance to discrete diffusion is challenging because guidance terms are not differentiable with respect to the discrete representations ${\mathbf{z}}_{t}$ (Gruver et al., 2024). Here, we introduce two guidance algorithms for discrete diffusion that circumvent the issue of non-differentiability. As in Section 2.2, we formalize the guidance term as a probability $p(y|{\mathbf{z}})\in\Delta^{K}$, where $y\in\{1,\ldots,K\}$ is one of $K$ possible classes, and we begin with the distribution that controls the strength of the guidance term via the $\gamma$ temperature parameter (with $c$ denoting the logarithm of the normalization constant):
$\log p^{\gamma}({\mathbf{z}}_{s}\mid y,{\mathbf{z}}_{t})=\gamma\log p(y\mid{\mathbf{z}}_{s},{\mathbf{z}}_{t})+\log p({\mathbf{z}}_{s}\mid{\mathbf{z}}_{t})+c.$ (5)
@@ 3.2 Classifier-based Guidance
Discrete Classifier-Based Guidance  We first introduce the following additional notation: given ${\mathbf{z}}^{(1:L)}$, let $\tilde{\mathcal{Z}}_{\ell}({\mathbf{z}}^{(1:L)})$ be the set of sequences $\tilde{{\mathbf{z}}}^{(1:L)}$ for which $\tilde{{\mathbf{z}}}^{(\ell^{\prime})}={\mathbf{z}}^{(\ell^{\prime})}$ for all $\ell^{\prime}\neq\ell,$ i.e., the set of sequences that are either the same as or only differ in position $\ell$ relative to ${\mathbf{z}}^{(1:L)}.$ We can sample from $p^{\gamma}({\mathbf{z}}_{s}^{(\ell)}\mid{\mathbf{z}}_{t}^{(1:L)},y)$ by training a classifier $p_{\phi}:\mathcal{V}^{L}\rightarrow\Delta^{K}$ on noised latents ${\mathbf{z}}_{t}^{(1:L)}$ for $t\in[0,1]$ and use this to model the first term on the right-hand side of (10) by only evaluating $p_{\phi}$ on sequences for which ${\mathbf{z}}_{s}^{(1:L)}$ and ${\mathbf{z}}_{t}^{(1:L)}$ differ by at most the token at position $\ell$:
@@
First Order Approximation  While tractable, this formulation suffers from the drawback that at each denoising step we must perform ${\mathcal{O}}(L\cdot N)$ forward passes through the classifier model, which can quickly become impractical for larger vocabularies and longer sequences. Similarly to Grathwohl et al. (2021), Vignac et al. (2022), we can treat the classifier $p_{\phi}$ as a continuous function of the one-hot inputs $\tilde{{\mathbf{z}}}^{(1:L)}\in\mathbb{R}^{L\times N}$ and use the first order Taylor approximation of $\log p_{\phi}$ to efficiently compute $p_{\phi}(y\mid\tilde{{\mathbf{z}}}^{(1:L)})$ with only a single forward and backward pass through the classifier model:
@@ 4 Uniform Diffusion Language Models
While masked diffusion models have demonstrated better language modeling compared to other discrete diffusion (Austin et al., 2021; Lou et al., 2023), we argue that they are less amenable to guidance, since once a token is unmasked at some time $t$ it remains so for all $s<t$. In contrast, for uniform noising, intermediate latents can be refined multiple times during the denoising process. We therefore revisit categorical uniform noise discrete diffusion, where $\bm{\pi}=\bm{u}:={\bm{1}}/N.$ Our aim is that by analyzing this class of diffusion models more carefully, we can reduce the gap to absorbing-state and yield performant models that are more easily steered by the guidance tools we developed above.
@@ 4.1 Uniform Noise Diffusion
Discrete Time Likelihood Bound  We start with the variational lower bound that is defined by a general diffusion process. For discrete-time diffusion, i.e., some finite steps $T$, we define $t(i)=(i+1)/T$ and $s(i)=i/T$ for $i$ in $0,\ldots,T.$ Denoting the Kullback-Leibler divergence as $D_{\mathrm{KL}}$, the denoising network $p_{\theta}$ is trained to minimize a variational upper bound (NELBO), which is given by:
$\mathbb{E}_{q}\Bigg{[}\underbrace{-\log p_{\theta}({\mathbf{x}}|{\mathbf{z}}_{t(0)})}_{\begin{array}[]{c}{\mathcal{L}_{\text{recons}}}\end{array}}+\underbrace{\sum_{i=1}^{T}D_{\mathrm{KL}}[q({\mathbf{z}}_{s(i)}|{\mathbf{z}}_{t(i)},{\mathbf{x}})\|p_{\theta}({\mathbf{z}}_{s(i)}|{\mathbf{z}}_{t(i)})]}_{\begin{array}[]{c}{\mathcal{L}_{\text{diffusion}}}\end{array}}\Bigg{]}+\underbrace{D_{\mathrm{KL}}[q({\mathbf{z}}_{t(T)}|{\mathbf{x}})\|p_{\theta}({\mathbf{z}}_{t(T)})]}_{\begin{array}[]{c}{\mathcal{L}_{\text{prior}}}\end{array}},$ (13)
When clear, we drop the explicit dependence of $t$ and $s$ on the discrete step $i$.
@@
$q({\mathbf{z}}_{s}\mid{\mathbf{z}}_{t},{\mathbf{x}})=\mathrm{Cat}\left({\mathbf{z}}_{s};\frac{N\alpha_{t}{\mathbf{z}}_{t}\odot{\mathbf{x}}+(\alpha_{t|s}-\alpha_{t}){\mathbf{z}}_{t}+(\alpha_{s}-\alpha_{t}){\mathbf{x}}+\frac{(\alpha_{s}-\alpha_{t})(1-\alpha_{s})}{N\alpha_{s}}{\bm{1}}}{N\alpha_{t}\langle{\mathbf{z}}_{t},{\mathbf{x}}\rangle+1-\alpha_{t}}\right)$ (14)
Denoising Process  The optimal form for the reverse diffusion process $p_{\theta}$ matches (14): in fact setting $p_{\theta}$ to (14) reduces the KL terms in (13) to zero. However, setting $p_{\theta}$ to exactly (14) is not possible because it cannot be a function ${\mathbf{x}}$ (which $p_{\theta}$ is generating). Therefore, we introduce a predictive model ${\mathbf{x}}_{\theta}({\mathbf{z}}_{t},t):\mathcal{V}\times[0,1]\rightarrow\Delta^{N}$ of the ‘clean’ data given a noisy latent ${\mathbf{z}}_{t}$ at time $t$. We use ${\mathbf{x}}_{\theta}$ to parameterize the denoising process as $p_{\theta}({\mathbf{z}}_{s}\mid{\mathbf{z}}_{t})=q({\mathbf{z}}_{s}\mid{\mathbf{z}}_{t},{\mathbf{x}}={\mathbf{x}}_{\theta})$, yielding:
@@
Note that this minimizes the ${\mathcal{L}_{\text{diffusion}}}$ term in (13) precisely when ${\mathbf{x}}_{\theta}={\mathbf{x}},$ as desired. To simplify notation, we omit below the explicit dependence of ${\mathbf{x}}_{\theta}$ on $t$.